# DoubleDiffusion: Combining Heat Diffusion with Denoising Diffusion for Generative Learning on 3D Meshes

### v3

@@ 5 Experiments
Figure 4: Qualitative comparison between the (a) MDF (baseline) and (b) DoubleDiffusion (ours) on the Stanford Bunny with $\sim$5k number of vertices.
@@
Table 1: Quantity comparison between MDF (baseline) and DoubleDiffusion (ours) with CelebA-HQ on Stanford Bunny (number of vertices $\sim$5k). With comparable MMD, our method surpasses MDF in terms of COV with a large margin. Meanwhile, our model is 4.7x smaller and 8.1x faster than MDF.
Figure 5: Our method can efficiently handle the mesh with large number of vertices manifold mesh. Here we show our result with (b) $\sim$25k vertices, and (c) $\sim$52k vertices (Stanford Bunny), with a qualitive comparison between (a) bunny with $\sim$5k vertices. The meshes with higher vertices are watertight remeshing with the off-the-shelf watertight manifold tool ManifoldPlus [16].
                   (a)                                                                                     (b)
5.1 Distribution Generation on Manifold
@@ 5.2.2 Qualitative results
Generative qualitive along the geometry. Figure 6 showcases qualitative results from our per-category texture generation experiment on the ShapeNetCore “chair” category. Each column represents a different chair model, where DoubleDiffusion generates textures that adapt to the unique geometric structure of each mesh. The leftmost images in each row show the untextured chair modelsFig. 6 demonstrate the qualitative result of texture generation on the different input chairs. The generated color on the mesh are smoothly varying following the chairs geometry, and the generated texture demonstrated a certain level of the understanding to the geometry.
In each example, the generated color on the mesh smoothly varies in alignment with the geometry of the chair, demonstrating the model’s capability to respect and adapt to surface details. Zoomed-in regions highlight the continuity and consistency of the generated textures, especially in curved and intricate parts of the mesh, such as the chair backs, armrests, and legs.
View Consistency. We present the generated results on ShapeNetCore models from five different views in Fig. 7. This multi-view presentation demonstrates DoubleDiffusion’s ability to achieve view consistency in signal generation directly on the mesh, without the need for 2D unrolling of the mesh. This consistency across views highlights the model’s capacity to respect the underlying mesh structure.
@@ 7 Conclusion
Supplementary Material
In this supplementary material, we provide additional implementation details in  Appendix A, theoretical analysis in  Appendix B and qualitative results in Appendix C.
Appendix A Implementation Details
To give more precise description to the framework in Fig. 3, we have include the diagram to the detail implementation for data preparation in Fig. S1 and heat diffusion block in Fig. S2.
A.1 Input data preparation
To process the mesh structure, we begin by computing the cotangent-weighted Laplacian matrix as discussed in Sec. 4.1. Using eigenvalue decomposition, we then derive its eigenvalues and eigenvectors. Additionally, we compute the mass matrix and the $X$- and $Y$-basis vectors for the tangent plane at each vertex. These tangent plane bases can be utilized to scale the heat-diffused features via tangent gradients, following the approach outlined in [31].
Figure S1: Detail implementation of the input data preparation, for both mesh structure data and the vertex color data.
The extraction of vertex colors is crucial to our experiments, as no existing 3D dataset provides hundreds or thousands of color data points defined on a single mesh structure. Therefore, we follow the baseline experiments [12] to generate the color data distribution using the Stanford Bunny mesh. This process involves employing the off-the-shelf 3D visualization tool PyVista [35] to obtain texture coordinates for the vertices of the Stanford Bunny mesh. With these texture coordinates, we perform UV mapping to transfer image data from the CelebA-HQ dataset, resulting in a mapped color vector assigned to the vertices of the mesh.
For the ShapeNetCore dataset, we utilize PyTorch3D [ravi2020pytorch3d] to load the texture atlases corresponding to shapes in the chair category. In PyTorch3D, textures are represented as atlases associated with the mesh faces. To obtain vertex colors from these face-based textures, we perform weighted interpolation. Specifically, for each vertex, we identify all the triangular faces that include it. The area of each triangle is computed, and the texture colors of these triangles are weighted proportionally to their respective areas relative to the total area of all triangles containing the vertex. This ensures accurate color interpolation at the vertex level.
A.2 Implementation of Heat Diffusion Block
The computation of heat-diffused features on the mesh is illustrated in Fig. S2. This process takes place within the heat diffusion block, corresponding to sub-block (c) in Fig. 3. Here, the noised vertex color features are first weighted by the mass matrix to account for the areas of the surrounding faces. The weighted features are then transformed into the spectral domain using the eigenvectors of the mesh structure. In this domain, the features are scaled by heat diffusion coefficients, enabling adaptive learning of geometric features on the mesh.
Figure S2: Implementation details of the heat diffusion block
Next, the heat-diffused features are projected back to the original feature domain and concatenated with the noised color features to produce an aggregated feature. This aggregated feature is then passed to the Timestep-Aware Per-Vertex MLP, as depicted in block (d) of Fig. 3, to perform probabilistic denoised diffusion.
Appendix B Theoretical Analysis
In this section, we would like to provide more detail to justify why and how Laplacian Beltrami operator can be used for learning on the manifold surface.
B.1 Integral on the manifold
The intension of using the Laplacian Beltrami operator on the manifold and the mesh is arise from the laplacian’s integration formulation on the manifold surface, as advised in the Mesh Laplacian Processing [34].
To analyze and process scalar functions defined on a manifold, the Laplace-Beltrami operator provides a framework for differential and integral operations, analogous to how convolutional kernels operate in the image domain. Since a manifold is locally Euclidean, Sörkine [34] demonstrated that points on the manifold can be represented by differential coordinates $\delta$, which are conceptually similar to $\Delta x$ in one-dimensional Euclidean space:
$\delta_{i}$ $=\bm{v}_{i}-\frac{1}{d_{i}}\sum_{j\in N(i)}\bm{v}_{j}$ (S1)
$=\frac{1}{d_{i}}\sum_{j\in N(i)}(\bm{v}_{i}-\bm{v}_{j}),$
where $d_{i}$ is a weighted factor representing the degree of the vertex i, which is the number of connection of vertex i. $N(i)$ represent the neighboring vertex $\bm{v}_{j}$ of the vertex i, it can also be a cotangent weights as a common practice in manifold mesh processing (Sec 3.2). It’s worth noting that, if we take differential coordinates to the continuous form,
[…]
Eq.(S1) becomes an integration of vertex $\bm{v}_{i}$ among the infinite small region within the surface curve $\gamma$ with perimeter $|\gamma|$, and $\bm{v}$ are the points that lies in this infinite small closed surface curve. And the $\textit{Curv}(\bm{v}_{i})$ in right hand side of Eq.(S2) represent the mean curvature [34] at vertex $\bm{v}_{i}$, and $\bm{n}_{i}$ is the surface normal.
These differential coordinates can be fastly computed through laplacian beltrami operator as it is a discretization version of the integration operator on the manifold, where the differential geometry on the mesh is given by:
$L=I-D^{-1}A,$ (S3)
where $I\in\mathcal{R}^{n\times n}$ is the identity matrix, $D\in\mathcal{R}^{n\times n}$ is the diagonal matrix with the entries $D_{ii}$ representing the number of the connections of the neighboring vertices (degree). $A\in\mathcal{R}^{n\times n}$ is the adjacency matrix, $A_{ij}=1$ when the edge exits between vertex i and j.
By incorporating geodesic distance and the area of adjacent triangles, the differential coordinates on the manifold (Eq. S1) with respect to a central vertex can be reformulated using cotangent weights as:
$\delta_{ij}=\sum_{j\in N(i)}\frac{\cot\alpha_{ij}+\cot\beta_{ji}}{2}(\bm{v}_{i}-\bm{v}_{j})$ (S4)
where $\alpha_{ij}$ and $\beta_{ij}$ are the angles opposite the edge connecting vertices $i$ and $j$ in the triangles adjacent to this edge, these are illustrated in Fig. 2(a). Thus, the use of the cotangent weighted Laplacian matrix in Eq. 2 and Eq. 3 upon the manifold surface gives as the geometric properties of the mesh structure.
Appendix C Experiment Results
In this section, we provides additional experiment results for per-category shape conditioned texture generation result on other categories, such as trains Fig. S3, pianos Fig. S4 and faucets Fig. S5.
Figure S3: Additional qualitative result on train meshes.
Figure S4: Additional qualitative result on piano meshes.
Figure S5: Additional qualitative result on faucet meshes.

### v5

@@ 5 Experiments
5.1 Distribution Generation on Manifold
Figure 5: Qualitative comparison between the (a) MDF (baseline) and (b) DoubleDiffusion (ours) on the Stanford Bunny with $\sim$5k number of vertices. Our method can efficiently handle the mesh with large number of vertices manifold mesh. Here we show our result with (c) $\sim$25k vertices, and (d) $\sim$52k vertices. The meshes with higher vertices are watertight remeshed with the off-the-shelf watertight manifold tool ManifoldPlus [11].
@@ 5.1 Distribution Generation on Manifold
Table 1: Quantity comparison between MDF (baseline) and DoubleDiffusion (ours) with CelebA-HQ on Stanford Bunny (number of vertices $\sim$5k). With comparable MMD, our method surpasses MDF in terms of COV with a large margin.Meanwhile, our model is 4.7x smaller in size than MDF and 8.1x faster for the inference, with only 0.14 second per denoising step.
@@ 5.2 Per-category Texture on ShapeNetCore
Figure 7: Generated results on different categories. Line 1-2 is the result on the train category, line 3-4 is the generated result for piano category, line 5-6 is the texture generation on the faucet category.
@@ 5.2.2 Qualitative results
Generative qualitative along the geometry. Figure 6 showcases qualitative results from our per-category texture generation experiment on the ShapeNetCore “chair” category. Each column represents a different chair model, where DoubleDiffusion generates textures that adapt to the unique geometric structure of each mesh. The leftmost images in each row show the untextured chair models. Fig. 6 demonstrate the qualitative result of texture generation on the different input chairs. The generated color on the mesh are smoothly varying following the chairs geometry, and the generated texture demonstrated a certain level of the understanding to the geometry.
In each example, the generated color on the mesh smoothly varies in alignment with the geometry of the chair, demonstrating the model’s capability to respect and adapt to surface details. Zoomed-in regions highlight the continuity and consistency of the generated textures, especially in curved and intricate parts of the mesh, such as the chair backs, armrests, and legs. More generation results on different category in is Fig 7, each category are trained independently.
View Consistency. We present the generated results on chair models from five different views in Fig. 8. This multi-view presentation demonstrates DoubleDiffusion’s ability to achieve view consistency in signal generation directly on the mesh, without the need for 2D unrolling of the mesh. This consistency across views highlights the model’s capacity to respect the underlying mesh structure.